### v1

Abstract
The accelerating loss of biodiversity presents critical challenges for ecological research and conservation strategies. The preservation of biodiversity is paramount for maintaining ecological balance and ensuring the sustainability of ecosystems. However, biodiversity faces numerous threats, including habitat loss, climate change, and the proliferation of invasive species. Addressing these and other ecology-related challenges, both at local and global scales, requires comprehensive monitoring, predictive and conservation planning capabilities. Artificial Intelligence (AI) Foundation Models (FMs) have gained significant momentum in numerous scientific domains by leveraging vast datasets to learn general-purpose representations adaptable to various downstream tasks. This paradigm holds immense promise for biodiversity conservation. In response, we introduce BioAnalyst, the first Foundation Model tailored for biodiversity analysis and conservation planning. BioAnalyst employs a transformer-based architecture, pre-trained on extensive multi-modal datasets encompassing species occurrence records, remote sensing indicators, climate and environmental variables. BioAnalyst is designed for adaptability, allowing for fine-tuning of a range of downstream tasks, such as species distribution modelling, habitat suitability assessments, invasive species detection, and population trend forecasting. We evaluate the model’s performance on two downstream use cases, demonstrating its generalisability compared to existing methods, particularly in data-scarce scenarios for two distinct use-cases, establishing a new accuracy baseline for ecological forecasting. By openly releasing BioAnalyst and its fine-tuning workflows to the scientific community, we aim to foster collaborative efforts in biodiversity modelling and advance AI-driven solutions to pressing ecological challenges.
keywords:
Foundation Model, Deep Learning, Representation Learning, Ecology, Biodiversity
1 Introduction
Figure 1: BioAnalyst is the first large-scale multi-modal model for biodiversity, trained on 20 years of spatiotemporal data modalities. The model ingests 10 distinct modalities, encoding and aligning them to latent ecological representations via the 3D Perceiver IO encoder. It then processes the latent space with the 3D Swin Transformer backbone and decodes it back to produce accurate spatiotemporal predictions. BioAnalyst shows strong performance in downstream tasks like (i)biotic, (ii) abiotic features prediction, (iii) long horizon prediction (12 timesteps = 1 year), both across space and time and (iv) is easily fine-tunable for any downstream task.
Biodiversity, encompassing the variety of all life forms on Earth, is fundamental to the stability and resilience of ecosystems. However, this rich diversity is under unprecedented threat due to numerous factors such as climate change [1], pollution [2, 3], habitat destruction, over-exploitation of natural resources [4], and the introduction of invasive species [5]. These pressures have led to significant declines in species populations and ecosystem degradation, posing critical risks to human well-being by compromising essential ecosystem services, such as clean air, water, and fertile soil [6].
Addressing these challenges requires predictive models to understand ecosystem dynamics and quantify the impacts of interventions. This, in turn, raises the overarching question of how to integrate such insights into decision-support frameworks for biodiversity conservation. Traditional methods often rely on static models, such as species distribution maps [7], which lack real-time updates and fail to capture rapid environmental changes. The fragmented nature of biodiversity data, dispersed across various sources and formats [8], hinders effective data harmonisation and integration. Additionally, ecological systems are inherently complex and less understood compared to engineered systems, making accurate modelling and prediction arduous tasks. Uncertainties and knowledge gaps persist, particularly in identifying and accounting for unknown variables and intricate inter-species interactions [9].
Recent advancements in AI and the development of Foundation Models offer promising avenues to overcome these challenges [10]. FMs, pre-trained on large-scale datasets primarily through self-supervision, have revolutionised fields such as natural language processing [11] and computer vision [12], demonstrating remarkable adaptability across diverse tasks. While geospatial foundation models are increasingly applied in ecological research, biodiversity modelling presents distinct challenges due to its reliance on unique data modalities, such as species occurrence records, trait databases, and fine-scale environmental covariates, for which specialised FMs have yet to be developed. The complexity and heterogeneity of ecological data, including species occurrence records, genetic sequences, remote sensing imagery, climate data, and environmental variables, pose significant challenges for integration and scalability. Moreover, the lack of standardised protocols for data collection and model development further complicates the creation of comprehensive AI tools in this domain [13].
In response to these challenges, we introduce BioAnalyst, a Foundation Model specifically designed for biodiversity analytics, opening new avenues on both local and global scale conservation planning efforts. Our contributions in this work are threefold:
  • Development of the first Multi-modal Biodiversity Foundation Model: We present BioAnalyst as the first large-scale AI model tailored for biodiversity modelling, capable of processing and integrating diverse data types to model complex ecological phenomena.
  • Advancement in Predictive Biodiversity Analytics: We demonstrate BioAnalyst’s predictive capabilities in key applications such as species distribution modelling, biotic and abiotic reconstruction, and population trend forecasting, especially in data-scarce scenarios.
  • Open Collaboration and Resource Sharing: By openly releasing BioAnalyst’s code, weights, and fine-tuning workflows, we aim to foster collaborative efforts within the scientific community, thereby accelerating research and conservation initiatives that address pressing ecological challenges.
2 Related Work
One of the first successful applications of FMs in Earth Sciences involves a geospatial foundation model trained on raw satellite imagery, called Prithvi, which can tackle tasks such as flood mapping, wildfire scar segmentation, multi-temporal crop segmentation, and cloud gap imputation [14]. A follow-up work introduced Prithvi WxC, a larger 2.2 billion-parameter FM that emulates weather and climate phenomena in tasks such as autoregressive rollout forecasting, downscaling, gravity wave parameterisation, and extreme events estimation [15]. On the same theme, Pangu-Weather delivered higher performance in medium-range forecasting, improving numeric weather prediction methods by training a 3D transformer model on 39 years of global data, which injects Earth-specific priors [16].
Focusing on Earth system dynamics and predictability, as well as the more specific and accurate prediction of extreme weather and climate events, ORBIT showcased advanced performance and highlighted the requirement for High Performance Computing (HPC) [17]. Similarly, Aurora can produce operational forecasts for global medium-range weather with unprecedented accuracy and speed-up over classical numerical weather prediction (NWP) models, by combining a flexible 3D Transformer backbone with a distinct encoder-decoder architecture [18]. Following similar approaches, Aardvak Weather features an end-to-end pipeline for data-driven weather prediction focusing on computation and maintenance benefits compared to classic NWP models [19]. In the Earth Observation domain, TerraMind is a large FM pre-trained on nine distinct modalities, highlighting the powerful alignment of token and pixel-level representations. In addition, this work demonstrates both the benefits of early fusion on downstream tasks and the performance gains achieved when learning on modalities generated by the FM [20]. To stimulate the development of FMs for earth monitoring, GEO-bench offers a suite of six classification and six segmentation tasks, suited for model evaluation [21].
In ecology, the number of FMs is relatively small, with a focus on visual, audio, and natural language tasks. More specifically, BioCLIP is an FM classifier for biology for the tree of life, trained on the TREEOFLIFE-10M, namely the abundance and variety of images of plants, animals, and fungi, together with the availability of rich structured biological knowledge [22]. Similarly, Insect-Foundation introduced a 1M dataset with insect imagery and taxonomy, and an FM based on ViT backbone [12] trained on this dataset for classification [23]. For species distribution modelling, NicheFlow demonstrated good predictive performance, mainly in reptile species [24], employing a Variational Autoencoder architecture and using environmental and species distribution variables. Combining audio and textual information NatureLM uses a pretrained encoder and a frozen LLM backbone (Llama 3.1-8b) to produce a text sequence used for bioacoustics tasks and more specific species-classification and detection [25].
3 Method
BioAnalyst has been designed to utilise the predictive power of the latest AI transformer-based models while being flexible enough to digest multi-modal geospatial input variables. Our work is inspired by the development of large-scale climate and weather models, such as Aurora [18] and Prithvi [15, 14], extending the capabilities of Foundation Models in the domain of biodiversity. More specifically, we are interested in learning about and forecasting biodiversity dynamics at both global and local scales with adequate resolution.
The design choices of BioAnalyst were driven by specific capabilities that it should possess, including multi-modal data representation, spatiotemporal feature preservation, global and local scale operation, underlying physics simulation across multiple scales, and various use-case adaptability. To account for them, BioAnalyst can be thought of as a forecast emulator, i.e., given a state of the Earth’s biodiversity at times $t$ and $t-\delta t$, it predicts the state at $t+\delta t$, where $\delta$ is a discrete time step. Although this might seem very simple, it poses significant challenges for modelling and engineering in complex domains such as ecology and, more specifically, biodiversity, which we have attempted to tackle to the best of our ability given the constrained resources at hand.
3.1 Foundation Model Architecture
Forecasting is a common task in Earth Sciences, such as weather, climate, and ecology, which is mainly modelled with time-series methods. Related work on weather and climate utilises the latest advancements in computer vision literature, including masked autoencoders [26], which exploit their low memory footprint, masking properties, and the handling of ungridded and sparse observation data.
BioAnalyst implements an encoder-backbone-decoder architecture. Let the input data at some time $t$ be a multi-modal tensors $\mathbf{X}_{t}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times C_{in}}$, representing $C_{in}$ variables over a spatial grid of height $\mathcal{H}$ and width $\mathcal{W}$. The model components are:
  • Encoder $\mathcal{E}$: We use Perceiver IO [27, 28], a general-purpose attention architecture. Input variables $\mathbf{X}_{t}$ are first tokenized into $N_{p}=\mathcal{H}/p\times\mathcal{W}/p$ non-overlapping patches of size $p\times p$. Fourier features encode spatial coordinates, which are combined with learned embeddings for variable types, time steps, and atmospheric levels. The resulting features associated with each patch are projected into the model’s embedding dimension $D_{e}$, creating tokens $\mathbf{T}_{t}\in\mathbb{R}^{N_{p}\times D_{e}}$. These are processed by the Perceiver IO encoder $\mathcal{E}$, which maps them to a fixed-size latent array $\mathbf{Z}_{t}\in\mathbb{R}^{N_{l}\times D_{e}}$ (where $N_{l}$ is the number of latent tokens) using cross-attention followed by self-attention layers:
    $\mathbf{Z}_{t}=\mathcal{E}(\mathbf{T}_{t})$ (1)
  • Backbone $\mathcal{B}$: We use a SwinTransformer [29] as the neural simulation engine. It receives the latent representations from two previous steps $\mathbf{Z}_{t-1},\mathbf{Z}_{t}$ and predicts the next latent state $\mathbf{Z}^{\prime}_{t+1}$ using hierarchical stages with shifted window self-attention:
    $\mathbf{Z}^{\prime}_{t+1}=\mathcal{B}(\mathbf{Z}_{t-1},\mathbf{Z}_{t})$ (2)
    This part aims to enable efficient computation while capturing spatial dependencies at various scales, thereby emulating the system dynamics in the latent space.
  • Decoder $\mathcal{D}$: The same Perceiver IO model is used to reconstruct the output variables. It makes use of specific query tensors $\mathbf{Q}\in\mathbb{R}^{N_{q}\times D_{e}}$ corresponding to the desired output variables (total $N_{q}$ features) and their coordinates on the target grid. These queries attend to the backbone’s output latent state $\mathbf{Z}^{\prime}_{t+1}$ via cross-attention within the decoder $\mathcal{D}$. The decoder outputs a sequence $\mathbf{\hat{Y}}_{t+1}\in\mathbb{R}^{N_{q}\times D_{e}}$ which is then projected and reshaped to the final multi-modal feature grid $\mathbf{\hat{X}_{t+1}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times C_{out}}}$ such that:
[…]
The design choices for BioAnalyst prioritise learning informative features in a compact latent representation before the emulation stage. By using Perceiver IO for both encoding and decoding stages, we aim to learn features from the original (raw) input data, thereby avoiding the standard approach of using separate tokenisation pipelines for each modality/variable-type, which can lead to biased tokens that are heavily dependent on the model used to produce them. This unified approach enables the model to capture cross-modal interactions at different data granularity levels, allowing it to differentiate features across various domains, from ground conditions and atmospheric levels to species distributions.
3.2 Pre-training Data Selection
BioAnalyst is pre-trained on BioCube [30], which compiles and aligns multiple datasets into a fixed spatio-temporal cube. Our main driver is modelling biodiversity ”as-a-whole”, which means we require observations from below the surface, the surface and above. Our analysis is confined to terrestrial (land-based) biodiversity; therefore, datasets describing marine or coastal biota are intentionally excluded from the study. More specifically, we select a subset of the total available features, categorised by modality groups, namely:
  • Atmospheric variables with 13 levels
  • Climate variables
  • Edaphic variables
  • Vegetation variables
  • Species Distribution variables
  • Land, Agriculture and Forest variables
  • RedList variables
  • Miscellaneous variables
These features are combined in a Data HyperCube, grounded on the coordinate reference system WGS84. The HyperCube contains global coordinates with a resolution of 0.25 degrees (grid sampling $\sim 28$ km) from the whole world while our focus is on European biodiversity, leading us to select a slice from it, yielding a Data Batch from [latitude: $\mathcal{H}$, longitude: $\mathcal{W}$ ] = [(32, 72), (-25, 45)] = [160, 280]. The observation time range spans from January 1, 2000, to June 1, 2020, and we sample with a 1-month lead time from this range.
Figure 2: A visual explanation of the data pipeline. From left to right, we received the data from BioCube in a HyperCube format, where sampling a single-timestep slice produces a Data Batch containing worldwide observations. Selecting European coordinates produces a Data Sample with multiple modalities stacked on the selected coordinate grid of size [160, 280].
Selecting a Data Sample from the Data Batch yields a composite multi-modal cell of European coordinates, with a specific monthly time-stamp. Each of these Data Points contains a total of 113 observations per location cell. More specifically, we denote the observed data points at a discrete time $t$ by a collection of tensors $\mathbf{X_{t}}$:
$\mathbf{X}_{t}=\sum_{i=1}^{10}\sum_{j=1}^{13}variables_{i}\cdot(levels_{j}),\mathbf{X}_{t}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times 113}$ (4)
A visual representation of the above can be found on Figure 2 while the complete data description is available in Appendix B.
3.3 Pre-training Objective
Pre-training climate and weather FMs for forecasting is frequently done by minimising a performance metric, either Mean Squared Error (MSE) or Mean Absolute Error (MAE). A straightforward approach is to force the model’s output at time $t+1$ to match the known future data. Formally, for a single variable $v$:
$\mathcal{L}_{MAE}=||\hat{x}_{t+1}^{v}-x_{t+1}^{v}||$ (5)
where $\hat{x}_{t+1}^{v}$ is the model’s prediction for a variable $v$ at time $t+1$. Summing across all variables and levels provides a multi-target objective.
In ecological contexts, temporal difference learning can be beneficial. Instead of predicting $x_{t+1}$ directly, we predict the increment $\Delta x_{t}=x_{t+1}-x_{t}$. This approach, often encountered in reinforcement learning [31], can reduce biases from unobserved global offsets or stable large-scale means. For instance, daily vegetation changes or seasonal fluctuations in species population can be smaller and more stable than absolute population numbers. By focusing on differences, we encourage the model to learn transition dynamics or, more specifically, biodiversity dynamics:
$\mathcal{L}_{TD}=||\hat{\Delta x_{t}^{v}}-(x_{t+1}^{v}-x_{t}^{v})||$ (6)
This choice is reinforced by empirical results in specific biodiversity modelling tasks (e.g., ephemeral wetlands or short-lived insect populations), which show improved forecasting stability over standard next-step MSE [32].
Going one step forward, given a system state $\mathbf{X}_{t}$ at time $t$, we aim to predict the next state $\mathbf{X}_{t^{\prime}}$ at time $t^{\prime}>t$. In the common single-step forecast scenario, following [18], we define a simulator function
$\Phi:(\mathbf{X}_{t-1},\mathbf{X}_{t})\rightarrow\mathbf{\hat{X}}_{t+1},$ (7)
which given two consecutive system states $\mathbf{X}_{t-1}$ and $\mathbf{X}_{t}$, predicts the future state $\mathbf{\hat{X}}_{t+1}$. Once we learn $\Phi$, we can roll out predictions over extended horizons in an autoregressive manner. Concretely, after setting $\mathbf{\hat{X}}_{t}=\mathbf{X}_{t}$ and $\mathbf{\hat{X}}_{t-1}=\mathbf{X}_{t-1}$, we can write:
$\mathbf{\hat{X}}_{t+k}=\Phi(\mathbf{\hat{X}}_{t+k+2},\mathbf{\hat{X}}_{t+k-1}),\text{for}\hskip 1.0ptk=1,2,...$ (8)
so, the next predicted state depends on the two most recent states, specifically the last real or predicted step. This repeated application of $\Phi$ is referred to as an iterative or autoregressive rollout.
3.4 Fine-tuning
BioAnalyst is fine-tuned in three different settings, each contributing to a distinct goal. In this section, a description of each setting is provided, and in the next section, we present the quantitative results.
3.4.1 Short-lead time finetuning
In this setting, we follow a similar approach to [18] and fine-tune the entire BioAnalyst for six rollout steps, effectively predicting biodiversity dynamics six months ahead. We freeze the whole architecture, including the encoder, decoder, and backbone, while training only the newly added VeRA adapters [33] on the backbone’s attention heads. We found VeRA to perform equally or sometimes slightly better than other Parameter Efficient Fine-tuning Techniques (PEFTs), such as LoRA [34], which use only one-tenth of the learnable parameters.
3.4.2 Roll-out finetuning
In this setting, following the same technique as before, we increase the trajectory length to 12 monthly observations, effectively predicting biodiversity dynamics one year in advance. The multi-step objective below is used for both short-lead and rollout fine-tuning settings.
$\mathcal{L}_{FT}=\frac{1}{K}\sum_{k=1}^{K}\mathcal{L}_{TD}(\hat{x}^{u}_{t+K},x^{u}_{t+K})$ (9)
The model is rolled forward $K$ time steps, where $K<horizon$, the loss is averaged across steps, and only the last step’s gradients are propagated due to the size of the model, following the push-forward trick [35].
3.4.3 Task specific finetuning
To evaluate the ecological capacity and environmental structure encoded in the BioAnalyst Foundation Model, we implement two complementary fine-tuning tasks. These tasks are designed to interrogate distinct dimensions of the model’s representation space: its ability to adapt to biotic presence-only data under temporal shift, and its capacity to retain structured abiotic gradients related to seasonal climate. For comparison, we also ran these fine-tuning tasks using the Aurora-0.25 model [18], which shares the same model architecture but has been pre-trained on climate modalities only.
The first task involves partial model adaptation: the BioAnalyst’s encoder and decoder are frozen. At the same time, the backbone is fine-tuned with VeRA adapters using historical species plant occurrence data from the GeoLifeCLEF2024 benchmark dataset [36] to forecast time-series distributions. The second task is diagnostic: a regression head is trained on top of frozen decoder embeddings for both BioAnalyst and Aurora, to predict monthly climate variables from CHELSA v2.1 [37]. Together, these tasks provide a dual lens on the model’s ecological generalisation and environmental coherence.
3.4.4 Biotic fine-tuning: forecasting species distributions
To assess whether BioAnalyst can learn to forecast biodiversity dynamics across time, we fine-tune the model’s backbone using anonymised plant species presence-only observations from the GeoLifeCLEF24 dataset. The model is trained on occurrences of 500 species (most frequent in the GeoLifeCLEF24 survey) across Europe from 2017 to 2020 and evaluated on its ability to predict species presence in 2021, without access to future climate or land-use data. The anonymity of the species focuses the modelling exercise on the process, rather than the species identity or phylogeny.
The model takes as input a species distribution matrix at time $t$ and is trained to predict the distribution at time $t+1$. The input and target distributions are normalised using species-specific statistics (mean and standard deviation). The model preserves the full spatial resolution of the data, working with grid-based representations rather than individual occurrence points. Training is performed end-to-end using a combination of loss functions, including the GeoLifeClef defined F1 score [38] and root mean square error (RMSE). The custom F1 score is determined in Equation 17. These metrics indicate whether the model learns both local and global distribution patterns, and also gives a comparable score for the GeoLifeCLEF benchmark.
BioAnalyst’s latent representations are then analysed through Principal Component Analysis (PCA) computation to understand how it encodes species-specific patterns. Visualising PCA helps assess whether the model learns meaningful environmental-species associations that can transfer across time points. This task simulates a realistic ecological forecasting scenario under observational uncertainty and evaluates whether the model’s environmental representation can be adapted to capture transferable species–environment associations.
3.4.5 Abiotic linear probing: recovering seasonal climate structure
To probe the fidelity of BioAnalyst’s pretrained environmental embeddings, we train a regression head to predict monthly climate values from the CHELSA v2.1 dataset against BioAnalyst’s decoder outputs. Specifically, we predict the monthly mean temperature ($\mathbf{tas}$) and total precipitation ($\mathbf{pr}$) over Europe ($0.25^{\circ}$ grids) for the years 2000 to 2019 (BioAnalyst is trained on data up to 2018). We reconstruct the decoder outputs for the same variables from BioAnalyst. For comparison, we used Aurora’s decoder predictions for 2-meter temperature. This creates a linear probing setup for evaluating how well the pretrained BioAnalyst latent representations encode climatically relevant information. By comparing the decoded reconstructions to the downsampled CHELSA targets, this linear probing setup assesses the alignment between learned representations and real-world climate signals, without updating the pretrained weights.
For each grid cell in the study area, the model predicts a 24-dimensional target vector representing the full annual cycle of temperature and precipitation (see Figure 3). We use mean squared error as the training objective and evaluate performance using RMSE, $R^{2}$, and monthly correlation metrics across diverse European bioclimatic zones. This probing task tests whether BioAnalyst’s pretrained latent space captures fine-grained abiotic structure, particularly seasonal variation, that is critical for ecological processes. As the encoder remains frozen, this setup isolates the representational quality of the pretrained embeddings, without the confounding effects of adaptation or fine-tuning.
Figure 3: Comparison between CHELSA target climate fields and decoded outputs from the BioAnalyst Small Decoder for May 2018 over Europe. Panels (a) and (b) show the downsampled CHELSA reference for precipitation (kg $m^{2}$) and temperature (K/10), respectively. Panels (c) and (d) show the corresponding decoded predictions from the model after reconstruction. The decoded precipitation captures major spatial gradients and orographic patterns (e.g., Alps, Norway), though with some smoothing. The decoded temperature field accurately reproduces latitudinal and coastal gradients present in the target. All maps are normalised for comparison.
4 Results
4.1 Rollouts: Forecasting biodiversity dynamics
In this part, we present four main results produced from our pre-training and rollout fine-tuning experiments. First, we present the cumulative mean for all the species distributions, comparing the ground truth distributions with the predicted distributions, considering only their absolute numerical values on Figure 4. The predictions closely follow the data trend, capturing both upward and downward changes in the species distributions which validates BioAnalyst’s capacity to effectively capture species distribution patterns.
Second, we report strong Sørensen similarity score of 0.31 for 28 species across all land grid cells which means that the predicted assemblage reproduces roughly one-third of the observed community patterns. Figure 5 highlights various patches that coincide with densely sampled or under-sampled regions. Overall, the spatial pattern confirms moderate compositional skill, useful for broad biogeographic inference, yet leaving room for improving species co-occurrence dynamics. Third, Figure 6 depicts BioAnalyst ability to localise and spatially predict granular species distributions.
Finally, we highlight the temporal drift of BioAnalyst’s species distribution predictions when performing 12 rollout steps on Figure 11. The species 1898286, 2491534, 8077224 and 9809229 exhibit the highest MAE while all the rest exhibit variable behaviour with the majority increasing the MAE on every next step.
Figure 4: The cumulative mean of all the species distributions in a 12 step rollout, predicting 1 year ahead.
Figure 5: The community Sørensen similarity map produced using Equation 19 with a mean value of 0.31 for the species variable group. Warm shades on the map highlight hotspots where more than half the species are matched, while cool blues mark cells with little or no overlap.
a) Refer to caption
b) Refer to caption
Figure 6: (a) Ground truth and prediction spatial plots for the species Pieris brassicae (ID 1920506) on 01-4-2019 and MAE of 0.00003. (b) Zoomed in plot, highlighting areas of interest, where the model can capture the general distribution, although failing to capture high-density areas.
4.2 Biotic fine-tuning: forecasting species distributions
To assess the biotic predictive capacity of BioAnalyst’s pre-trained embeddings, we fine-tuned a classification head to forecast species presence in the GeoLifeCLEF 2021 dataset. The model achieved high accuracy with an F1 score of 0.9964 and an RMSE of 0.5284, demonstrating its ability to support fine-scale species distribution modelling. In the same setting, Aurora performs marginally worse in evaluation metrics than BioAnalyst, as shown in Table 1. Nevertheless, investigating the predicted species richness in more detail for both models, we found that Aurora overestimates its predictions, while BioAnalyst is more spatially accurate, as shown in Figure 7. This is an expected result, as BioAnalyst has been trained on 28 species distributions and exhibits better spatial grounding.
Figure 7: Comparison of observed and predicted species richness in the Netherlands for the year 2021. a) The Observed Richness based on empirical field data from the GeoLifeCLEF2024 survey. b) The species richness for 2021 predicted by the BioAnalyst, and c) the predictions from the Aurora Model for 2021. The results have been masked for land only, as both models predicted richness for most ocean pixels as well.
We further analysed the learned representation structure via Principal Component Analysis (PCA) on the backbone outputs. The first three principal components accounted for 96.4% of the total variance, with PC1 alone explaining 54.44%, followed by PC2 (31.05%) and PC3 (10.92%) (Figure 8). This concentration of variance in a few dimensions suggests that the BioAnalyst latent space captures most biotic information in a compact and low-dimensional manifold, facilitating robust downstream generalisation across ecological prediction tasks.
Figure 8: Correlation matrix of the first six principal components (PCs) derived from the BioAnalyst backbone in the GeoLifeCLEF task. The near-zero off-diagonal values indicate that the principal components are orthogonal, confirming effective dimensional separation in the learned latent space.
Table 1: Performance of the species distribution forecasting task predicting species distribution for 2021.
Model Loss F1 RMSE
BioAnalyst 0.0057 0.9964 0.5284
Aurora 0.0130 0.9945 0.5014
4.3 Abiotic linear probing: recovering seasonal climate structure
The model achieved strong predictive performance, with the best epoch reaching an $R^{2}$ of 0.9002, a loss of 0.0225, and an RMSE of 0.1499, as shown in Table 2. These values indicate that the BioAnalyst decoder outputs contain sufficient information to reconstruct fine-grained seasonal climate patterns even without any fine-tuning of the encoder. In the same theme, comparing BioAnalyst with the pre-trained Aurora-025 yielded a stronger predictive capacity for BioAnalyst. However, this comparison is not entirely representative, as it provides a performance baseline, since BioAnalyst is trained at a monthly frequency, while Aurora is trained at a 6-hour interval.
To further examine the structure of the learned embeddings, we applied Principal Component Analysis (PCA) to the decoder outputs. The resulting correlation matrix of the first six principal components revealed strong orthogonality between components, as expected Figure 9. This suggests that the BioAnalyst representations organise abiotic variation along separable axes of climate seasonality and geography, supporting their use in downstream ecological tasks.
Figure 9: Monthly aggregated correlation matrix of the first six principal components (PCs) derived from the BioAnalyst decoder outputs in the CHELSA linear probing task from 2000-2019. The near-zero off-diagonal values indicate that the principal components are orthogonal, confirming effective dimensional separation in the learned latent space. This supports the interoperability of downstream regression analyses along independent axes of climate variability.
Table 2: Performance of the linear probing task predicting down-sampled CHELSA v2.1 temperature and precipitation targets from BioAnalyst decoder outputs over Europe.
Model Loss R2 RMSE
BioAnalyst 0.0225 0.9002 0.1499
Aurora 0.2668 0.7354 0.5144
5 Conclusion & Discussion
In this work, we introduced BioAnalyst, the first Foundation Model for biodiversity. BioAnalyst is truly multi-modal, light-weight and can be used to model various complex ecological phenomena with competitive performance, setting a new accuracy baseline for ecological forecasting. We highlighted its predictive analytics as a stand-alone model in tasks such as biodiversity and dynamics modelling, as well as in tasks like species distribution forecasting, absence detection, and monthly climate linear probing.
The results demonstrate that Foundation Models pre-trained on heterogeneous ecological and environmental data can generalise well across a range of predictive biodiversity tasks. Notably, BioAnalyst scales to regional or continental domains without incurring the computational costs typically associated with larger geospatial models. The model’s strong performance in both biotic and abiotic tasks suggests that it learns a unified latent representation of environmental structure that is transferable and biologically meaningful. This opens new opportunities for rapid model adaptation in data-poor contexts and for advancing hypothesis-driven ecological modelling through representation learning. In addition, we have utilised open, available, and licensed data. We have also open-sourced the model weights, training routine, scaling recipe and fine-tuning techniques used during our experiments. We believe that the research community can be greatly benefited by using our pre-trained weights but also inspired by our open-sourced solution to pursue new avenues of research in complex fields with multi-modal data, complex modelling and scalability demands.
The avenues for improvement are numerous, as BioAnalyst has set the proving ground for the application of Foundation Models in ecology, specifically in biodiversity. Future extensions may include incorporating additional modalities, such as genomic or functional trait data, to enable even more expressive and interpretable biodiversity analytics. At present, BioAnalyst produces exclusively deterministic forecasts. Adopting a probabilistic framework is very important for variables that behave stochastically, like species geographic distributions, surface latent-heat flux, precipitation, convective snowfall rate, 10m wind, and more [39]. Towards that end, a promising line of future work could investigate ensembles of BioAnalyst models trained on different datasets or fine-tuning the model into a probabilistic variant.
Another promising avenue for improvement would be the generation of biodiversity-related modalities that stem from EO FMs, such as TerraMind, and continue to fine-tune them for new tasks. It would be interesting to see the correlations between these modalities and the raw modalities used in the pre-training of BioAnalyst. Additionally, expanding BioAnalyst’s capabilities with user interaction features, such as chat, would enhance its interpretability.
However, some caution is warranted when interpreting downstream predictions. For example, while BioAnalyst correctly captures trends such as species distribution decline in the data (Figure 4), this may partly reflect temporal biases in observation effort rather than actual ecological change. Such artefacts highlight the need for careful disentanglement of signal and sampling in biodiversity datasets when training and evaluating foundation models. Future work could address this by incorporating observation effort metadata or by explicitly modelling detection processes alongside ecological predictors.
A key open question for any AI-based modelling is the quantification of uncertainty in the model’s predictions. Adding extra dimensions, such as spatial and temporal, makes the task even more challenging. Methods for quantifying this uncertainty at different granularity levels, such as cartograms [40] or meta-model traits [41], are part of BioAnalyst’s future work. Truly synergistic models that embed ecological principles (e.g. energy budgets, trophic interactions) into AI architectures that simulate population dynamics are an exciting frontier [42]. This could mean neural networks that respect mass-balance constraints or reinforcement learning agents that simulate animal foraging behaviour. Such integration would yield models that not only predict well but also adhere to known ecological laws, making them more generalisable, trustworthy, and aligned with global biodiversity goals [43].
Like any other advanced AI model, BioAnalyst has limitations. One of the biggest is its constrained area of operation, Europe, which is not representative of global biodiversity dynamics. Additionally, BioAnalyst is trained on data that represent biodiversity only at the terrestrial level. It does not take into consideration the sea, for example, amplifying the bias towards other parts of biodiversity and the compound effects they may have on both global and local biodiversity dynamics. Higher-quality, more frequently sampled, and better-curated data points could further enhance BioAnalyst’s capabilities in combination with longer training. As mentioned earlier, the uncertainty quantification of BioAnalyst’s prediction is a project in itself and was not the focus of this work.
Declarations
[…]
Supplementary information
A supplementary file accompanies this article.
[…]
Acknowledgements
This study has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No 101057437 (BioDT project, https://doi.org/10.3030/101057437). Views and opinions expressed are those of the author(s) only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the European Commission can be held responsible for them.
This publication is part of the project Biodiversity Foundation Model of the research programme Computing Time on National Computing Facilities that is (co-) funded by the Dutch Research Council (NWO). We acknowledge NWO for providing access to Snellius, hosted by SURF, through the Computing Time on National Computer Facilities call for proposals. This work utilised the Dutch national e-infrastructure, supported by the SURF Cooperative, under grant no. EINF-10148.
[…]
Author contributions
A.T. Core contributor, conceptualised and formulated the research, original draft, implemented and evaluated all aspects of BioAnalyst, including the model, the architecture, the training and fine-tuning routines, designed all the experiments, evaluations and visualisations, while managing the project M.M. Main contributor, implemented and evaluated all aspects of BioAnalyst, including the model, the training and fine-tuning routines, all the experiments and evaluations S.S. Main contributor, engineering the data pipeline S.G. Main contributor, architecting of the model T.K. Main contributor, originated and carried out the biotic and abiotic fine-tuning tasks, prepared the task datasets models, and visualisations, provided feedback on the evaluation D.P. Set up and managed the distributed infrastructure of the project, carried long-time experiments A.v.V. Supervised the project and provided guidance and feedback All authors contributed to the writing and editing of this manuscript.
[…]
Code availability
All our code is open-sourced under the MIT license and can be found at:
[…]
Data availability
The dataset used can be found at https://huggingface.co/datasets/BioDT/BioCube and its code base at https://github.com/BioDT/bfm-data.
Appendix A BioAnalyst Implementation Details
This appendix provides a detailed description of the BioAnalyst model’s architecture, as a supplement to section 3.1.
A.1 Core Architectural Blueprint
The BioAnalyst model is structured as an encoder-backbone-decoder system. Let the input data at time $t$ be a multi-modal tensor $\mathbf{X}_{t}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times C_{in}}$ , where $\mathcal{H}$ and $\mathcal{W}$ represent the spatial height and width of the input grid, and $C_{in}$ denotes the number of input variables (channels).
Firstly, $\mathbf{X}_{t}$ is first processed by an encoder $\mathcal{E}$ module built upon Perceiver IO, which transforms the input data into a fixed-size latent representation $\mathbf{Z}_{t}\in\mathbb{R}^{N_{l}\times D_{e}}$, where $N_{l}$ is the number of latent tokens and $D_{e}$ is the embedding dimensions. The encoder can process inputs from one or more time steps (e.g., $t$ and $t-1$) to form $\mathbf{Z}_{t}$. This stage includes methods, including positional and temporal encodings, as detailed in later sub-appendices.
Next, $\mathbf{Z}_{t}$ is then fed into a backbone network $\mathcal{B}$. BioAnalyst muses a Swin Transformer as its backbone. The Swin Transformer processes $\mathbf{Z}_{t}$ through hierarchical stages with shifted window self-attention to model spatio-temporal dynamics and predict $\mathbf{Z}^{\prime}_{t+1}$.
Finally, $\mathbf{Z}^{\prime}_{t+1}$ is passed to a decoder $\mathcal{D}$, which uses a set of learnable query vectors $\mathbf{Q}$ corresponding to the desired output variables and their target grid locations, which attend to $\mathbf{Z}^{\prime}_{t+1}$ to reconstruct the multi-modal feature grid $\mathbf{\hat{X}}_{t+1}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times C_{out}}$.
A.2 The BioAnalyst Encoder
The encoder $\mathcal{E}$ transforms the raw, multi-modal input tensor $\mathbf{X}_{t}$ into a structured, fixed-size latent representation $\mathbf{Z}_{t}$.
A.2.1 Input Processing and Feature Engineering
The spatial dimensions ($\mathcal{H}$ and $\mathcal{W}$) of each variable in $\mathbf{X}_{t}$ are first divided into non-overlapping patches. Each patch is of size $p\times p$ (where we kept $p=4$ in BioAnalyst’s configuration). This results in $N_{p}=(\mathcal{H}/p)\times(\mathcal{W}/p)$ patches per variable. For each variable type, the data within each patch potentially spanning multiple channels (e.g., different variables within a group), is flattened and then linearly projected to form initial patch tokens. To provide rich contextual information, these tokens are combined with several learned embeddings:
  • Spatial coordinate encoding: the normalized centroid coordinates $(x,y)$ (i.e., $x,y\in[-1,1]$)) for each patch are encoded using Fourier features. For $N_{f}$ frequency bands (which in our case were set to $N_{f}=64$ and a maximum frequency of 224), the features for each coordinate are $[\sin(s_{0}\pi x),\cos(s_{0}\pi x),\dots,\sin(s_{N_{f}-1}\pi x),\cos(s_{N_{f}-1}\pi x)]$, where $s_{k}$ are linearly spaced frequencies. These are concatenated with the original coordinates and projected to the model’s embedding dimension $D_{e}$.
  • Variable-specific embeddings: distinct embedding layers are used for different categories of variables (e.g., surface, atmospheric, species) to distinguish their semantic meanings. These include dedicated embeddings for different atmospheric pressure levels and individual species channels.
  • Temporal embeddings: both the absolute timestamp of the input and the forecast lead time $\delta t$ are encoded using sinusoidal functions and projected through separate linear layers.
The initial patch tokens are then combined by summing them with these various embeddings. The resulting feature-rich tokens $\mathbf{T}_{t}\in\mathbb{R}^{N_{total}\times D_{e}}$ (where $N_{total}$ is the total number of tokens generated across all patches and variable types/levels), form theinput to the Perceiver IO’s attention mechanisms.
A.2.2 Perceiver IO Latent Transformation
The encoder maps the input tokens $\mathbf{T}_{t}$ to a fixed-size latent array $\mathbf{Z}_{t}\in\mathbb{R}^{N_{l}\times D_{e}}$ using a two-stage process. First, a set of $N_{l}$ learnable latent query vectors distill information from the input tokens via cross-attention:
$\text{CrossAttn}(\mathbf{Q}_{lat},\mathbf{K}_{T},\mathbf{V}_{T})=\text{softmax}\left(\frac{\mathbf{Q}_{lat}\mathbf{K}_{T}^{\top}}{\sqrt{d_{k}}}\right)\mathbf{V}_{T}$ (10)
where $\mathbf{Q}_{lat}$ are derived from the learnable queries, and $\mathbf{K}_{T},\mathbf{V}_{T}$ are the keys and values derived from the input tokens $\mathbf{T}_{t}$, and $d_{k}$ is the dimension of the keys. The resulting latent array is then processed through a stack of self-attention layers (a Transformer tower) to allow latent tokens to interact and refine the representation. To manage computational load, this module employs Grouped-Query Attention (GQA) and standard regularization techniques like Layer Normalization and Dropout.
A.3 The BioAnalyst Backbone
The backbone ($\mathcal{B}$) serves as the neural simulation engine, taking the latent representation $\mathbf{Z}_{t}$ from the encoder and predicting the state for the next time step, $\mathbf{Z}^{\prime}_{t+1}$. It was designed as 3D Swin Transformer architecture, structured as a U-Net. This design includes an encoder path that progressively downsamples the latent representation and a decoder path that symmetrically upsamples it, with skip connections linking corresponding stages to preserve high-resolution details.
A.3.1 Backbone Encoder Path
The Swin Transformer backbone takes in $\mathbf{Z}_{t}$, a latent tensor including temporal information. $\mathbf{Z}_{t}$ consists of $N_{l}$ tokens arranged in a 3D latent grid ($N_{ld}\times N_{lh}\times N_{lw}$), with each token holding $D_{e}$ features.
The tensor passes through multiple encoder stages, each made of Swin Transformer blocks that apply self-attention across latent features. After each stage (except the deepest one – the “bottleneck”), patch merging halves spatial dimensions (depth, height, width) and doubles feature dimensionality, enabling coarser feature extraction.
The output features from each encoder stage, specifically the features before the path merging operation, are preserved. The preserved features are then passed to the corresponding stages in the decoder path via skip connections.
A.3.2 Backbone Decoder Path
The decoder starts from the bottleneck features and mirrors the encoder structure. Each decoder stage (except the first) begins with patch splitting, which doubles spatial resolution and halves feature dimensionality, preapring features for fusion with encoder outputs.
Skip connections combine upsampled decoder features with matching encoder outputs via element-wise addition, except at the highest resolution, where concatenation is used. This concatenated output is linearly projected to restore the feature dimension $D_{e}$.
Each stage then applies Swin Transformer blocks. The final decoder output, $\mathbf{Z}’_{t+1}$, matches the input $\mathbf{Z}_{t}$ in shape ($N_{l}\times D_{e}$) and represents the predicted next latent state.
A.3.3 Hierarchical Processing with Shifted Windows
As it could be noticed, the core computational unit within both the encoder and deocoder paths of the U-Net backbone is the Swin Transformer block. The Swin Transformer processes latent volumes through a series of these blocks. The number of blocks per stage and the number of the attention heads per block are configurable hyperparameters.
The following characteristics of the Swin Transformer blocks can be considered:
  • Windowed Multi-Head Self-Attention (W-MSA): self-attention is computed within local 3D windows (e.g., of size $W_{D}\times W_{H}\times W_{W}$, where $W_{D},W_{H},W_{W}$ are window dimensions for latent depth, latent height, and latent width respectively). This reduces computation compared to global self-attention, as attention is restricted to non-overlapping local windows.
  • Shifted Window Multi-Head Self-Attention (SW-MSA): to allow for cross-window connections, consecutive Swin Transformer blocks, alternate between regular W-MSA and SW-MSA. In SW-MSA, the window configuration is shifted by half a window size relative to the previous layer. This cyclic shift ensures that the creation of a larger receptive field over layers.
  • Relative position bias: relative position biases are added to the attention scores, potentially improving generalization across different locations within the windows.
  • Multi Layer Perceptron (MLP) layers: each attention module is followed by a 2-layer MLP with Gaussian Error Linear Unit (GELU).
A.4 The BioAnalyst Decoder
The BioAnalyst decoder ($\mathcal{D}$) translates the predicted latent state $\mathbf{Z}^{\prime}_{t+1}$ from the backbone back into a high-resolution, multi-modal grid of observable variables $\mathbf{\hat{X}}_{t+1}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times C_{out}}$. Similar to the encoder, the decoder uses a Perceiver IO core, allowing for flexible and targeted generation of outputs.
A.4.1 Output Query Formulation
The decoder uses a set of specific, learnable output queries, each targeting a specific variable at a given spatial location and time. For each target variable map (e.g., surface temperature or species distribution), a query vector $\mathbf{q}\in\mathbb{R}^{D_{e}}$ is formed by combining:
  • Target variable embedding: identifies the variable to predict (e.g., ’surface temp’, ’species X’).
  • Spatial coordinate embedding: spatial positions over the $(\mathcal{H},\mathcal{W})$ grid are encoded via Fourier features, projected to $D_{e}$, and interpolated to match the number of output queries ($N_{q}$), yielding a shared spatial tensor of shape $N_{q}\times D_{e}$.
  • Temporal embedding: encodes the forecast time step $t+1$ using a lead time embedding and optionally, an absolute time encoding.
  • Atmospheric level/Species index embedding (if applicable): distinguishes pressure levels or species indices for level-specific or species-specific outputs.
These components are summed to form each query vector. Collectively, the queries form a query array $\mathbf{Q}\in\mathbb{R}^{N_{q}\times D_{e}}$ (where $N_{q}$ is the total number of distinct output variable maps to be generated), covering all defined outputs (surface, single-level, atmospheric, species, land etc.).
After attending to $\mathbf{Z}^{\prime}_{t+1}$, the decoder produces output embeddings of shape $N_{q}\times D_{e}$. These are projected through task-specific layers into flat variable maps, then reshaped into ×$\mathcal{H}\times\mathcal{W}$ grids—one for each target variable, level, or species.
A.4.2 Cross-Attention with Backbone Output
The output queries $\mathbf{Q}$ attend to the final latent state $\mathbf{Z}^{\prime}_{t+1}$ produced by the Swin Transformer backbone. This cross-attention mechanism allows each query to selectively extract the relevant information from the dense latent representation needed to predict its specific target. The attention operation is analogous to that in the encoder:
$\mathbf{\hat{Y}}_{t+1}=\text{CrossAttn}(\mathbf{Q},\mathbf{K}_{Z^{\prime}},\mathbf{V}_{Z^{\prime}})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}_{Z^{\prime}}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V}_{Z^{\prime}}$ (11)
where $\mathbf{K}_{Z^{\prime}}$ and $\mathbf{V}_{Z^{\prime}}$ are keys and values derived from the backbone’s output $\mathbf{Z}^{\prime}_{t+1}$. The result, $\mathbf{\hat{Y}}_{t+1}\in\mathbb{R}^{N_{q}\times D_{e}}$, is an array where each row corresponds to an output query and contains the decoded information in the embeddings dimension.
Essentially, the Perceiver IO architecture used for the decoder focuses on the cross-attention between the output queries and the backbone’s output latent state.
A.4.3 Projection and Reshaping to Final Output
The Perceiver IO decoder outputs a sequence of embeddings, $\mathbf{\hat{Y}}_{t+1}$, where each embedding must be projected to the actual value of its target variable. A variable-specific linear projection maps each $D_{e}$-dimensional embedding to the appropriate output shape. For scalar outputs (e.g., temperature at a location), this means projecting to a single value.
Dedicated projection layers handle each variable group, converting embeddings into structured outputs. For example, atmospheric variables are projected into tensors shaped by batch size, number of variables, number of pressure levels, height, and width. Species-related outputs include an extra species dimension.
After projection, all outputs are reshaped and organized into the final multi-modal grid $\mathbf{\hat{X}}_{t+1}\in\mathbb{R}^{\mathcal{H}\times\mathcal{W}\times C_{out}}$, where $C_{out}$ is the total number of predicted variable channels.
A.5 Positional, Temporal, and Variable-Specific Encodings
Effective representation of spatial, temporal, and categorical information is highly-important for BioAnalyst to interpret inputs and generate accurate, context-aware predictions. This sub-appendix provides more details into the specific encoding schemes that were used.
A.5.1 Temporal Encoding Schemes
BioAnalyst encodes two temporal aspects: the absolute time of an observation and the forecast lead time.
  • Absolute time encoding: The calendar date and time of an input (or decoder target) time step $t$ are converted to a scalar value $\tau$, then encoded using sinusoidal functions. For embedding dimension $D_{e}$, the resulting vector $\mathbf{e}_{time}\in\mathbb{R}^{D_{e}}$ has components: $(\mathbf{e}_{time})_{2i}=\sin(\tau/10000^{2i/D_{e}})$ and $(\mathbf{e}_{time})_{2i+1}=\cos(\tau/10000^{2i/D_{e}})$ for $i\in[0,D_{e}/2-1]$. This encoded vector is then processed by a learned linear layer to match $D_{e}$ and allow for learnable adaptation.
  • Lead time encoding: the forecast lead time, $\Delta t=t_{forecast}-t_{input}$, is also encoded.This scalar value (e.g., 2 months) undergoes the same sinusoidal encoding as above and is then projected using a separate learned linear layer. This informs the model of how far into the future it is forecasting.
Both the encoder and decoder components use these time encodings, adding them to the patch or query embeddings to provide temporal context.
A.5.2 Variable-Specific and Categorical Feature Embeddings
To differentiate between the various input modalities and their specific characteristics, BioAnalyst uses learned embeddings for different categories of data:
  • Variable type embeddings: each input variable (e.g., 2m temperature, species extinction risk) gets a unique, learnable embedding of size $D_{e}$. Separate linear layers are used for different variable groups, projecting the patchified data (which implicitly includes variable identity due to how data is batched and fed) into the shared embedding space.
  • Atmospheric level embeddings: for atmospheric variables variables across pressure levels (e.g., 50 hPa, 500 hPa), each level’s spatial data is tokenized and projected using a shared linear layer. These level-specific tokens are then concatenated with others, letting the model differentiate vertical context via token position.
  • Species index embeddings: similarly to atmospheric levels, for multi-species variables, each species channel is tokenized and projected via a shared layer. These species-specific tokens are concatenated to enable attention layers to learn species-aware features.
All embeddings are learned during training and added to the patch tokens before being fed into Perceiver IO, providing spatial, temporal, and semantic context for forecasting.
A.6 Data Normalization
BioAnalyst normalizes all variables before processing them in the encoder and unnormalizes the outputs of the decoder to produce the final predictions. All the variables are normalized separately, and the variables which have more levels are normalized per-level (e.g., species distributions and atmospheric variables). For the normalization and denormalization, we compute statistics across the whole dataset by collecting mean values and standard deviations. The relationship between normalized and unnormalized variables is the following:
$\mathbf{X}^{t}_{v,i,j,\text{normalized}}=\frac{\mathbf{X}^{t}_{v,i,j}-\text{centre}_{v}}{\text{scale}_{v}}$ (12)
Appendix B Dataset
This section provides a detailed description of the data used to pre-train and finetune BioAnalyst.
B.1 Introduction
BioAnalyst is pre-trained on a part of BioCube as discussed at section 3.2, using a Batch structure throughout with statistics available at Table 5.
B.2 Variable groups
For our data mixture we have used the below variable groups $V$ with their corresponding variables $v$:
  • Surface variables: t2m, msl, slt, z, u10, v10, lsm
  • Edaphic variables: swvl1, swvl2, stl1, stl2
  • Atmospheric variables: z, t, u, v, q
  • Pressure levels: 1000, 925, 850, 700, 600, 500, 400, 300, 250, 200, 150, 100, 50
  • Climate variables: smlt, tp, csfr, avg_sdswrf, avg_snswrf, avg_snlwrf, avg_tprate, avg_sdswrfcs, sd, t2m, d2m
  • Miscellaneous variables: avg_slhtf, avg_pevr
  • Vegetation variables: NDVI
  • Land variables: Land (percentage of total land area)
  • Agriculture variables: Agriculture, Arable, Cropland (percentage of land area)
  • Redlist variables: RLI (indicator of the changing state of global biodiversity)
  • Forest variables: Forest (percentage of land area)
  • Species variables: Species occurrences records
This variables selection is inspired and extracted from various works around species distribution modelling, habitat assesement and ecosystem modelling with classical and Machine Learning methods.
More specific, climatic energy and moisture variables like temperature and humidity reflect thermal nichies where humidity gives vapour-pressure deficit a key plant and anthropogenic stressor [44]. Precipitation indicates the water balance which is the strongest global predictor after temperature [45]. Radiation on various lengths (short and long) drives photosynthesis and evapotranspiration and has showed improvement in Net Primary Productivity (NPP)-linked biodiversity models [46]. Snow variables at high latitudes and mountains control growing season length and habitat suitability.
For edaphic water and temperature status, we selected to look into soil moisture which is a direct proxy for plant water stress and root-zone dynamics and has a strong interactive effect with species richness [47]. Soil temperature provides germination cues linked with soil microbial activity while variables like potential evapotranspiration and surface latent heat flux are need for water deficit estimation [48].
For vegetation, we select Normalized Difference Vegetation Index (NDVI) that quantifies the health and density of vegetation and is a strong indicator of the greenness of the biomass. Additionally, we select a series of indicators that quantify the monthly changes of land, agriculture (arable and cropland).
Finally, for species we select 2 variables to work with. The first is the Red List Index (RLI) which is an indicator of the changing state of global biodiversity and defines the conservation status of major species group and measures trends in extinction risk over time. The second is the species occurrences records for 6 major categories that is highly ranked in importance from European Union. More specific Large carnivores, farmland birds, wild pollinators, herpetofauna, invasive alien species (IAS) and Mediterranean endemics together span the core biodiversity priorities currently driving European nature policy and funding. Large carnivore management has become a political flash-point: the Nature Restoration Regulation explicitly references “conflict species” such as the wolf, and the European Parliament voted in May 2025 to downgrade the wolf from strict to general protection under the Habitats Directive [49]. Farmland birds remain the EU’s headline biodiversity indicator; the European Environment Agency reports a $40\%$ decline in the Farmland Bird Index since 1990, underscoring the need for landscape-scale restoration [50]. Wild pollinators stand at the centre of the revised EU Pollinators Initiative (“A new deal for pollinators”), which commits all Member States to continent-wide monitoring and trend reversal by 2030 [51]. IAS require surveillance and early-warning systems under Regulation (EU) 1143/2014, which obliges Member States to establish national monitoring and rapid-response mechanisms [52]. Finally, herpetofauna and Mediterranean endemics receive targeted LIFE funding and are focal taxa in Article 17 conservation-status reporting under the Habitats Directive, anchoring them in the EU’s mandatory assessment cycle. Table 3 lists in detail all the species from the above categories that are used for training BioAnalyst.
Table 3: Approximate total occurrences between 2000-2020 and species ID to scientific name mappings according to GBIF for six major species categories
Major Category Species ID Scientific Name (Common) Total occurrences
Farmland birds 8077224 Alauda arvensis (skylark) 2.5 M
2491534 Emberiza citrinella (yellowhammer) 2.5 M
2473958 Perdix perdix (grey partridge) 400 k
4408498 Crex crex (corncrake) 140 k
9809229 Sturnus vulgaris (common starling) 5.0 M
Herptiles 2431885 Triturus cristatus (great crested newt) 149 k
8909809 Emys orbicularis (European pond turtle) 39 k
2430567 Pelobates fuscus (spadefoot toad) 12 k
Invasive & Alien Species 8002952 Ambrosia artemisiifolia (common ragweed) 87 k
2437394 Callosciurus erythraeus (Pallas’s squirrel) 900 k
3034825 Heracleum mantegazzianum (giant hogweed) 181 k
2891770 Impatiens glandulifera (Himalayan balsam) 422 k
5218786 Procyon lotor (raccoon) 36 k
Large Carnivores 5219173 Canis lupus (grey wolf) 22.5 k
2433433 Ursus arctos (brown bear) 5.8 k
2435240 Lynx lynx (Eurasian lynx) 34.9 k
5219219 Canis aureus (golden jackal) 1.4 k
5219073 Gulo gulo (wolverine) 5.7 k
Mediterranean Species 2435261 Lynx pardinus (Iberian lynx) 436 k
5844449 Aquila fasciata (Bonelli’s eagle) 55 k
2441454 Testudo hermanni (Hermann’s tortoise) 34 k
2434779 Monachus monachus (Mediterranean monk seal) 137 k
8894817 Caretta caretta (loggerhead sea turtle) 6.5 k
Pollinators 1340503 Bombus terrestris (buff-tailed bumblebee) 266 k
1340361 Bombus hyperboreus (Arctic bumblebee) 325 k
1898286 Vanessa atalanta (red admiral) 2.0 M
1920506 Pieris brassicae (large white) 1.8 M
1536449 Episyrphus balteatus (marmalade hoverfly) 0.01 k
Table 4: ERA5 variable names, their definitions and measure units [53] [54]
chat Variable Units Description
swvl1 m3 m–3 Volumetric soil water content, layer 1 (0cm depth)
swvl2 m3 m–3 Volumetric soil water content, layer 2 (7cm depth)
stl1 K Soil temperature, layer 1 (0cm depth)
stl2 K Soil temperature, layer 2 (7cm depth)
smlt m (water equivalent) Snow melt accumulated at surface
tp m Total precipitation (liquid + frozen)
csfr kg m–2 s–1 Convective snowfall rate
avg_sdswrf W m–2 Mean surface downwelling shortwave radiation flux
avg_snswrf W m–2 Mean surface net shortwave radiation flux
avg_snlwrf W m–2 Mean surface net longwave radiation flux
avg_tprate m s–1 Mean total precipitation rate
avg_sdswrfcs W m–2 Mean surface downwelling shortwave flux (clear sky)
sd m Snow depth
t2m K 2m air temperature
d2m K 2m dew point temperature
msl Pa Mean sea level pressure
slt code Soil type classification code
z m2 s–2 Geopotential
t K Air temperature (pressure levels)
u m s–1 Eastward wind component (pressure levels)
v m s–1 Northward wind component (pressure levels)
u10 m s–1 10m eastward wind component
v10 m s–1 10m northward wind component
q kg kg–1 Specific humidity (pressure levels)
lsm 0/1 Land-sea mask (0=sea, 1=land)
avg_slhtf W m–2 Mean surface latent heat flux
avg_pevr kg m–2 s–1 Mean potential evaporation rate
B.3 Build Batches
We have developed a pipeline of systematic data batch construction to enable learning across diverse and temporally aligned environmental and biodiversity signals. Each batch combines multiple data modalities over a fixed temporal window of one calendar month, including both the beginning and end time points. Its design ensures compatibility with geospatial deep learning models that require structured tensors across space, time, and modality. The pipeline creates batches with a temporal resolution of two consecutive monthly slices (e.g., January and February 2001). Spatially, it follows a fixed grid at $0.25^{\circ}\times 0.25^{\circ}$ resolution, matching the Copernicus ERA5 grid and covering the full anticipated range of longitudes and latitudes within Europe. All variables are re-projected via appropriate transformations to this common spatial resolution to ensure alignment across modalities. The pipeline supports the data sources and modality groups introduced in Section B.2.
The batching process is fully deterministic: for a given input dataset and time window, it produces outputs without random variability. This design enables consistent benchmarking across runs and supports long-term model evaluation on fixed test splits. Furthermore, the use of non-overlapping monthly intervals ensures temporal independence between batches, which is essential for forecasting and change detection tasks.
Every input dataset is preprocessed and harmonized based on the following principles:
  • Longitude coordinates are wrapped into the interval $(-180^{\circ},180^{\circ}]$ to ensure consistency across global datasets.
  • Timestamps are standardized to datetime64 objects with monthly resolution.
  • Latitude and longitude coordinates are snapped to the $0.25^{\circ}$ grid to match the spatial resolution of the batch format.
  • Missing values are imputed with zeros, enabling compatibility with models that do not natively support NaN values.
To maintain modularity and extensibility, the pipeline separates data loading logic per modality (e.g., _load_era5, _load_csv, _load_species), with all outputs standardized into spatially and temporally aligned xarray.Dataset objects. This design supports the seamless addition of future data types such as Sentinel-2 imagery, elevation, or anthropogenic indicators without altering the core batch assembly logic.
The batching pipeline is optimized to handle large-scale datasets efficiently. ERA5 NetCDF files are processed using xarray with chunking enabled, and CSV files are filtered and indexed spatially using precomputed $0.25^{\circ}$ grid maps. This architecture allows scaling to continental datasets and long temporal ranges without excessive memory usage.
Xarray reads ERA5 NetCDF files, combines them by coordinates, and retains exactly two calendar months per batch. If multiple temporal slices exist within the same month, only the earliest one is kept to reduce redundancy. Land, agriculture, or vegetation CSV files are ingested and parsed into month-specific rasters. The pipeline supports both common CSV structures:
  • Layout A: includes a Variable column and individual year columns (e.g., 2000, 2005).
  • Layout B: includes variable-year columns directly (e.g., NDVI_2020, Land_2015).
If expected variables or time slices are missing in a particular file, the system logs the missing entries and fills them with zeros. This ensures that tensor dimensions remain consistent across batches, even when data coverage is sparse for certain modalities or regions.
Species data are extracted from Parquet files. Each species is treated as an independent raster tensor per month, based on the reported Distribution value. A master species list is inferred from the available data Table 3 to ensure consistent dimensionality across batches. This design enables both single-species and multi-species modelling and supports spatially explicit predictions of species presence.
All variables are converted into PyTorch tensors. Scalar geospatial variables (e.g., temperature, NDVI) are stored in tensors of shape $(2,H,W)$, where $H$ and $W$ are the grid height and width and 2 accounts for the consecutive timestamps. Pressure-level variables (e.g., geopotential, wind at altitude) are stored as $(2,C,H,W)$ where $C$ is the number of pressure levels. Species presence tensors are organized as one tensor per species with shape $(2,H,W)$. In cases where pressure-level variables are included, their corresponding pressure levels are saved in the batch metadata.
Each batch includes a metadata dictionary capturing the timestamp window, grid specification (latitude and longitude), list of included species, and pressure levels (if applicable). The final batch is a structured dictionary with modality keys representing the various variable categories (e.g., surface_variables, agriculture_variables, species_variables), each containing variable-specific tensors. The complete batch is serialized and saved to disk in PyTorch’s binary .pt format, enabling efficient loading during model training without repeating and preprocessing operation.
Table 5: Statistics of the constructed data Batches used in the pre-training of BioAnalyst
Attribute Value
Grid Sampling Resolution 0.25°
Global Grid Size 1440 (lon) $\times$ 720 (lat) = 1,036,800 cells
Europe Grid Size 280 (lon) $\times$ 160 (lat) = 44,800 cells
Latitude Bounds [32.0°, 72.0°]
Longitude Bounds [–25.0°, 45.0°]
Time Range 01–01–2000 to 01–06–2020
Number of Batches 233
Batch Size (average) $\sim$43 MB
Total Storage Volume 10 GB
Temporal Resolution per Batch 2 months
Pressure Levels Included 13
Species per Batch 28
Total Data Points 5.062.400
B.4 Fine-tune datasets
For task-specific finetuning we have used two different datasets for the corresponding tasks respectively. More specific we have used:
GeoLifeCLEF24: GeoLifeCLEF is benchmark dataset with a training set of close to 5 million plant occurrences in Europe (single-label, presence-only data) as well as a validation set of about 5000 plots and a test set with 20000 plots, with all the present species (multi-label, presence-absence data) [36]. A visual depiction of how the data look like can be found on Figure 10.
Figure 10: Yearly aggregated plant species richness per $0.25^{\circ}\times 0.25^{\circ}$ grid across Europe (35 – 60 ° N, 10 ° W – 30 ° E) for the GeoLifeCLEF24 survey data from 2018-2021 with 5 million occurrences. Lighter colors show higher abundance.
CHELSA v2.1: CHELSA (Climatologies at High resolution for the Earth’s Land Surface Areas) delivers global grids of mean, minimum and maximum 2 m air temperature (tas, tasmin, tasmax) and precipitation rate (pr) at 30-arc-second resolution ($\sim 1$ km) from 1979 - present. Temperature fields are produced through lapse-rate–based statistical downscaling of ERA-5/ERA-Interim reanalyses, while precipitation is downscaled with an orographic algorithm that incorporates wind fields, boundary-layer height and subsequent bias-correction with Global Precipitation Climatology Centre (GPCC) gauge data, yielding markedly improved representation of mountain rain-shadows and thermal gradients compared with coarser products [37]. These high-fidelity temperature and precipitation layers form the foundation for CHELSA’s 19 bioclimatic derivatives and for time-series forcings such as CHELSA-W5E5, and are now a de-facto standard in species-distribution modelling, trait–environment analyses, hydrological and vegetation-dynamics models, and climate-change impact assessments. Because ecological responses to climate are often threshold-driven and spatially heterogeneous, the $\sim 1$ km detail of CHELSA allows modellers to resolve local refugia, elevational turnover and fine-scale moisture stress that coarser climatologies fail to capture, thereby increasing predictive accuracy and reducing uncertainty across taxa and regions [55].
Appendix C Implementation details
This appendix provides detailed information about the model card, the hyperparameters the training recipe, the software used and the methods used for efficient scaling, as a supplement to sections 3.3 and 3.4
C.1 Model card
We have implemented and star training 2 versions of BioAnalyst, one Small sized with 440M parameters and one Medium with 980M parameters. The complete model configuration can be found on Table 6 and Table 7. The idea behind these configurations is the increasing and decreasing sizes of the kernel dimensions, following our U-Net style backbone architecture.
Table 6: Model card.
Model patch_size num_heads embed_dim depth swin_backbone_size Model Size
Small 4 12 384 6 medium 440M
Medium 2 16 512 10 large 980M
Table 7: Swin‐backbone configurations for Large and Medium model sizes
Parameter Large Medium
encoder_depths [2, 2, 2] [2, 2]
encoder_num_heads [8, 16, 32] [8, 16]
decoder_depths [2, 2, 2] [2, 2]
decoder_num_heads [32, 16, 8] [16, 8]
window_size [1, 4, 5] [1, 1, 1]
mlp_ratio 4.0 4.0
qkv_bias True True
drop_rate 0.0 0.0
attn_drop_rate 0.0 0.0
drop_path_rate 0.1 0.1
Furthermore, to enhance performance and computational efficiency, particularly given the high dimensionality of the input and output data, several techniques are used. Group-Query Attention [56] is used within the Perceiver IO attention layers to reduce the memory footprint associated with key-value projections during cross-attention. Additionally, standard regularization techniques such as Layer Normalization and Dropout are applied throughout the Perceiver modules. The Swin Transformer backbone makes use of stochastic depth [57] to improve generalization by randomly skipping the residual branch during training, thus reducing the depth on a per-sample basis.
C.2 Software
For the development of BioAnalyst we used the Python programming language [58] and developed the required modules on PyTorch’s neural network library [59] and PyTorch Lightning [60]. For visualisation we employed the Streamlit package [61]. For data operations we used xarray package [62].
C.3 Training and Scaling
BioAnalyst has been trained for 1000 epochs - 80.000 gradient steps with a batch size of 1, using the AdamW optimiser [63] with cosine-annealing learning rate schedule with periodic warm restarts [64]. We used a starting learning rate of $0.00005$ and weight decay of $0.000005$ with $T_{periodic}=8000$ gradient steps.
Variable weighting
To balance the loss during pre-training and subsequent finetuning tasks we assign individual weights to each variable for every variable group. Our weighting scheme is inspired by the works of [18, 65, 66, 67, 68, 69, 70] and are reported on Table 8.
Table 8: Variable-specific weights $(w_{v})$ used in the loss function.
Variable Group $(V)$ Variable $(v)$ Weight $(w_{v})$
Surface t2m 2.50
msl 1.50
slt 0.80
z 1.00
u10 0.77
v10 0.66
lsm 1.20
Edaphic swvl1 1.10
swvl2 0.90
stl1 0.70
stl2 0.60
Atmospheric (p-levels) z_pl 2.80
t_pl 1.70
u_pl 0.87
v_pl 0.60
q_pl 0.78
Climate smlt 1.00
tp 2.20
csfr 0.60
avg_sdswrf 0.90
avg_snswrf 0.70
avg_snlwrf 0.50
avg_tprate 2.00
avg_sdswrfcs 0.50
sd 0.90
t2m_clim 2.50
d2m 1.30
Vegetation NDVI 0.80
Land cover Land 0.60
Agriculture Agriculture 0.40
Arable 0.30
Cropland 0.40
Forest Forest 1.20
Redlist RLI 1.30
Miscellaneous avg_slhtf 1.20
avg_pevr 1.00
Species species 10.00
Training Loss
As discussed in section 3.3 our pre-training objective is the temporal difference error $\mathcal{L}_{TD}$ which is
$\mathcal{L}_{TD}=||\hat{\Delta x_{t}^{v}}-(x_{t+1}^{v}-x_{t}^{v})||$ (13)
In addition, during training we weight each variable’s loss with the weights we defined before at Table 8
$\mathcal{L}_{\mathrm{TD}}(t)\;=\;\sum_{v\in\mathcal{V}}w_{v}\;\bigl{\|}\hat{\Delta x_{t}^{\,v}}-\bigl{(}x_{t+1}^{\,v}-x_{t}^{\,v}\bigr{)}\bigr{\|}_{1},$ (14)
Disclaimer
The results on this version of the manuscript, are produced with the Small model, while we strive to finalise training the Medium model.
C.4 Metrics
C.4.1 Pre-training
Mean Absolute Error (MAE): To evaluate the performance of our model during pre-training, we measure and log the MAE between the predictions and the ground truth target which is
$\text{MAE}=\frac{1}{V}\sum_{v=1}^{V}\frac{1}{H\times W}\sum_{i=1}^{H}\sum_{j=1}^{W}||\mathbf{\hat{X}}_{i,j}^{v}-\mathbf{X}_{i,j}^{v}||$ (15)
where $i,j$ index over the longitude and latitude dimensions $H,W$ of each variable $v\in V$.
C.4.2 Task-specific finetuning
Root Mean Square Error (RMSE): To evaluate the performance of the finetuned model in both biotic, abiotic tasks, we measure and log the RMSE between the predictions and the ground truth target which is
$\text{RMSE}=\frac{1}{V}\sum_{v=1}^{V}\sqrt{\frac{1}{H\times W}\sum_{i=1}^{H}\sum_{j=1}^{W}(\mathbf{\hat{X}}_{i,j}^{v}-\mathbf{X}_{i,j}^{v})^{2}}$ (16)
$\textbf{F}_{1}$ score: To evaluate the performance of biotic fine-tuning task and obtain a comparison metric with the downstream dataset used.
$F_{1}\;=\;\frac{1}{N}\sum_{i=1}^{N}\frac{\mathrm{TP}_{i}}{\displaystyle\mathrm{TP}_{i}+\frac{\mathrm{FP}_{i}+\mathrm{FN}_{i}}{2}}$ (17)
[…]
Coefficient of determination: To evaluate the performance of the abiotic linear probing task
[…]
where $y_{i}$ are the observed values, $\hat{y}_{i}$ are the predicted values, $\bar{y}$ is the sample mean of the observations and $n$ is the number of data points.
Sørensen similarity map: To compare the species sets in each grid-cell, using the incidence (presence/absence) form of the Sørensen–Dice coefficient
$S_{ij}\;=\;\frac{2\,c_{ij}}{2\,c_{ij}+b_{ij}+c_{ij}},$ (19)
where $c_{ij}$ is the number of species present in both ground-truth and prediction at cell $(i,j),b_{ij}$ those present only in the observation, and $c_{ij}$ only in the prediction. The mean assemblage similarity estimate is an informative indicator for the performance of the species distribution model (SDM) [71].
Appendix D Further Results
Figure 11: A scorecard highlighting the MAE from the autoregressive rollout (12 steps) for the species variable group